1. A random sample of 10 independent healthy people showed the following body temperatures (in degrees Fahrenheit) and summary statistics: Temps 98.50 98.20 99.00 96.30 98.30 98.70 97.20 99.10 98.70 97.20 Average 98.12 Std Dev 0.92

Compute a 95% CI for the average body temperature of healthy people.
Answer #1

Given that, D = 10 X = 98.12 s = 0.92 .id.l.: 0-1 = 10-129 At 95%. Con Ridence level the tis. x=1-o.g5 = 0.05 Y2 = 0.05 - 0.0